### v2

@@
[…]
Khai PhanFaculty of Applied Science,
University of British Columbia,
2332 Main Mall,
Vancouver, BC V6T 1Z4, Canada
[…]
University of British Columbia,
2207 Main Mall
Vancouver, BC V6T 1Z4, Canada
email: siddharth.rout@ubc.ca
Acoustics-based Active Control of Unsteady Flow Dynamics using Reinforcement Learning Driven Synthetic Jets
[…]
Résumé
This study proposes the use of deep reinforcement learning (DRL) to actively control wakes and noise from flow past a cylinder by leveraging acoustic-based pressure feedback. A hydrophone array captures downstream signals, enabling a DRL agent to adjust jet actuators on the cylinder’s surface in real-time. The method reduces noise levels by up to 9.5% and drag coefficients by up to 23.8%, effectively minimizing flow-induced vibrations. This highlights the potential of DRL-driven active flow control for engineering applications.
keywords:
Acoustic, Deep Q-Network, Drag Reduction, Reinforcement Learning, Wake Control
1 Introduction
@@
Flow control has always been one of the most anticipated engineering problems due to its ubiquitous applicability. From the suppression of flow oscillation in open cavities [1] to the construction of hybrid rocket motors [2], flow control has been used as an indicator of how technology has developed to counter the stochasticity of nature. Throughout the last few decades, the number of paradigms for flow control has been increasing more and more. Applications of flow control to air vehicle systems, including fixed wing airfoils, turbomachinery, combustion, aeroacoustics, vehicle propulsion integration, and rotorcraft. Flow control methods can be categorized into Active Flow Control (AFC) and Passive Flow Control (PFC). There are many innovative applications established in various industries[3, 4]. PFC have a constant control law that is consistent with time and do not get any feedback on how well the controller performs, such as having changes to aerodynamic shapes or textures. The passive methods include Gurney flap, vortex generator, bump, cavity, roughness, small disturbance, bleed, splitter plate, polymer, and biomimetic techniques[5, 6]. Some examples are leading-edge serrations, riblets, corrugated airfoils and lubricated skins. Most of these are widely implemented in aircrafts to delay flow separation and increase lift to drag ratio. Winglets are nowadays found to be used to reduce tip vortex formation to reduce drag. However, these control strategies are limited as the control can not be manipulated temporally based on feedback or requirements. For instance, what if the PFCs act adversely? So, AFC is a good way out as it can take in feedback from the state and actuate the controller intelligently. The active methods include oscillation and flow perturbation, acoustic excitation, jet, synthetic jet, plasma actuator, and Lorentz force. Many interesting AFCs developed in past decades. For example, installation of the synthetic jet to change the vortex shedding pattern[7]; utilizing wavelength actuators to attenuate turbulence[8]; and studying the effects of acoustic excitation on vortex shedding [9]. Among those numerous methods, the application of blowing-suction of velocity jets stands out as one of the most practical and widely recognized, evidenced by NASA’s experiment on the Boeing 757 with jet actuators incorporated in the vertical stabilizer to reduce drag and improve the overall performance of the plane [10, 11].
@@
Why specifically controlling flow past a cylinder? Flow-induced forces play determinant roles in the life and safety of structures as well. Oscillations in the flow cause fatigue, enhance defects, aeroelastic flutters, and decrease the factor of safety of structures. Falling of the famous Tacoma Narrows Bridge is a popular case of structural failure due to similar causes[12]. Tall buildings like Taipei 101, Burj Khalifa etc. have to be designed to be able to face fast winds[13, 14]. Passive techniques developed in the past few decades are still very promising due to the ease of utilization in industry[15, 16, 17]. AFCs past cylinders are excellent toy problems to demonstrate concepts. The oppositely placed suction and blowing around cylinder became popular active control strategy in 2000s[18, 19, 20]. To improve the quality nonlinear control algorithms[21] and eigensystem-realization based reduced order model for suppression of wakes[22] are popular. However, these deterministic control algorithms like proportional–integral–derivative(PID) controllers often require approximation of state space and calculation of the transfer function to actuate AFCs is expensive yet inaccurate and non-generalizable as the transfer function is case dependant. Hence, data-driven model-free methods like reinforcement learning (RL) for AFC are well appreciated as they are generalizable.
@@
In the past few years, there has been a surge in Deep Reinforcement Learning (DRL) based flow control techniques[23, 24, 25, 26]. This work is hence based on a DRL algorithm to be able to extend the work to realistic cases like flow past marine vehicles. A few recent attempts to utilize DRL in AFC include controlling two synthetic jets of blowing/suction [27] and implementing adjoint-based partial derivative equation augmentation to DRL to solve flow simulation more efficiently[28]. Both showed great results in controling wakes. In order to reduce vortex shedding, the two aforementioned papers both tried to decrease the drag coefficient in the simulation. However, no work is registered in which acoustic-based flow control model is used to reduce wakes formation and control flow dynamics.
@@
Deep Q-Nework (DQN) is a branch of DRL that involves calculating the Q values of each step the model takes and, much like other ML algorithms, learning to maximize the returned reward from the environment. DQN seems to function well in more abstract tasks and therefore is widely used in robotics and has achieved great success in this field. For instance, Fernandez-Fernandez et al. study the application of DQN to human-like sketching performed by robots[29]. In the context of of controls and optimization in fluid dynamics morphing airfoils and shape optimization using DQN is also evident[30]. Despite the level of sophistication of the model, it is not very widely used in AFC modelling. In addition, manipulation of other properties apart from drag calculation is not common in the overall scientific conversation in AFC. According to Klapwijk et al., turbulence in the fluid flows is the source of sound generation in the system. The article explores noise levels when the turbulent flow is increased[31]. Interestingly, it also claims that the noise generation mechanism is difficult to understand. In this work, it is shown that by controling vortex generated noise using deep Q-learning drag and wake amplitudes could be controlled.
@@
In the nineteenth century, wakes are identified as the major source of noise in the flow past an object. There lies the concept behind this work and it is theoretically supported. Strouhal in 1878 and Kohlrausch in 1881 independently found out about a faint sound originating from vortices, to which the latter described as ’reibungstone’ [32, 33, 34]. Sir James Lighthill, in 1950s, discovered the theoretical connection between fluid flow and acoustics from the conservation laws to derive the wave equation for acoustics, called the theory Lighthill’s Acoustic Analogy [35]. Sir Lighthill creates an experiment assuming a patch of turbulent flowing fluid surrounded by a large domain of surrounding stationary fluid. Let turbulent flow produce noise; however, the noise would transmit to the surrounding fluid at rest. By analysing and comparing the terms in the conservation equations for stationary fluid, the resulting equation could be written as a forced bidirectional wave equation. It becomes clear that the forcing term or the source term in the wave equation is what generates noise in flow. Here is the derived wave equation in Einstein’s notations,
$\frac{\partial^{2}\rho}{\partial^{2}t}-c_{o}^{2}{\nabla^{2}\rho}=\frac{\partial^{2}{T_{ij}}}{{\partial x_{i}}{\partial x_{j}}}$
${T_{ij}}={\rho}{v_{i}v_{j}}-{\sigma_{ij}}+(p-c_{o}^{2}\rho){\delta_{ij}}$
@@
where T is called Lighthill’s turbulence stress tensor and has three components or three sources for noise generation. ${\rho}{v_{i}v_{j}}$ is the convection of momentum fluctuation, $\mathbf{\sigma}$ is the viscous stress tensor and $(p-c_{o}^{2}\rho)\delta_{ij}$ is the difference in exact pressure $p$ and approximated thermodynamic pressure, $c_{o}^{2}\rho$. ${\rho}$ is the density, $c_{o}$ is the speed of sound, $t$ is the time dimension, x is the spatial dimension and v is the velocity. This equation quantifies sound sources from flowing fluid where it takes care of thermodynamics jumps, turbulent fluctuations, and viscous dissipation. Now the fact is if the stress tensor is non-zero sound is formed and with the origin of wakes, vortex stretching starts into action and even with laminar vortex street sound is produced. This sound is often considered tonal while a broadband noise is generated in case of turbulent flows. Various further works have demonstrated this [36, 37, 38].
@@
In this paper, due to the limitation of computational resources, the experiments are conducted using incompressible flow. This is based on this simple assumption that incompressible flows still produce acoustic-like pressure fluctuations, though incompressible flow solvers assume constant density. This is because the Lighthill’s stress tensor (T) is non-zero. The speed of sound in such cases tends to infinity. As such speed of sound being much slower than the speed of fluid, Mach number << 1, the nearfield observations are barely affected [39, 40, 41, 42]. We assume these pressure fluctuations as approximate noise. So, wakes make noise like pressure fluctuation and hence definitely if louder is the noise in a flow then stronger are the wake vorticities. Though this logic is very much valid but it has not been used for control or analysis rather flow states like pressure and velocity fields are considered as direct measurements and sometimes vorticity field is used as a direct measure of rotational energy in a flow. In this research, we aim to minimize the wake formation in the flow and hence lower the specially calculated effective sound pressure levels (SPLs) created by the vortices in the flow past a stationary circular cylinder using the flow-generated sound. Source of sound is a better and easier way to control vorticity in a flow as vorticity is the source of generated noise, hence being a more logical measure for such flow control problems. Furthermore, we will explore the effect of lowering SPLs on the oscillating drag experienced by the cylinder. As for the active control algorithm aspect of the research, DQN based reinforcement learning is used to control the blowing-suction of two synthetic jets on opposite ends of the cylinder perpendicular to flow.
@@
2 Deep Q-Learning
Constructed on the Markov Decision Process in which the quality of action at a particular state is learned based on the reward due to the action, Q-Learning, has been a very popular early reinforcement learning[43]. Conceptually, the action at a particular state is independent of the historical state following Markov probability. However, the rewards are learned from the cumulated score of rewards in an episode of the control process. The convergence of the optimality control problem using the Bellman Equation under stochastic updates was proved soon after[44]. The limitation of Q-Learning is the finite nature of the map between state to best action, which is called Q-table. Deep neural networks as excellent maps in Q-Learning replace the Q-table to get called Deep Q-Network (DQN) and the algorithm is called Deep Q-Learning. It is a breakthrough reinforcement learning algorithm since it has been able to match human-level control of console to Atari games[45]. DQN allows mapping to conditional non-linear control algorithms. The second benefit is the possibility of being trained in an infinite and continuous environment state space.
Figure 1: Interaction between DQN agent and Environment in Markov Decision Process
@@ 2 Deep Q-Learning
$\mathbf{V}(\mathbf{s})=max(\mathbf{R}(\mathbf{s},\mathbf{a})+\gamma\mathbf{V}(\mathbf{s^{\prime}}))$
@@
Figure 2: Diagram of flow past cylinder setup
@@
Regarding the construction of the neural network of the DQN Agent, four fully connected layers are made between the input and output layers. The input layer takes in the calculated SPL and the output layer gives out the two jet velocities. Each layer consists of 50 neural nodes which let 7902 learnable parameters. ReLU is the nonlinear activation function used at each node, which essentially acts as switch, hence making DQN a multilayer nonlinear switch. Adam optimizer[46] in PyTorch library is applied to maximize rewards returned in each episode. More information about the utilization of the DQN agent in different tasks will be provided in the following section. The purpose of this task is to minimize the sound pressure levels (SPL) created by the wakes past the cylinder. There are multiple specific setup details in order to obtain the SPL reduction.
3 Problem introduction and setup
The computational setup is built on DOLFINx, which is a high-performance solver of partial differential equations written in C++ for backend integration with legacy FEniCSx(version 2019.1.0) and python for interface[47, 48]. The project allows the use of the standard benchmark case "Flow past a cylinder (DFG 2D-3 benchmark)”, as a simulation framework to further develop the research based on [49, 50, 51]. The setup includes a horizontal rectangle with a height of 0.41m and a length of 2.2m, and the bottom left corner of the rectangle is at coordinate (0, 0). The obstacle is a circular-based cylinder with a radius of 0.05m centered at coordinate (0.2, 0.2). As the flow develops its oscillation, though laminar, the obstacle will experience a drag force, $C_{D}$, which can be determined using the formula:
$C_{D}=\frac{2}{\rho U^{2}_{mean}L}\int_{\partial\Omega_{S}}\{\rho\nu\mathbf{n.}\mathbf{\nabla}u_{t_{s}}(t){n_{y}}-p(t){n_{x}}\}ds$

### v3

@@
Abstract
Flow generated noise are caused shear flows and, hence, they can be used as feedback to control the flow. Existing flow control uses state variables like velocity, pressure, or vorticity, none use acoustic observables as the primary control signal. It is tough to model a classical control algorithm using sound level but data-driven approaches are not as do not have to explicitly model the physics. We present an acoustics-driven framework for active control of unsteady wake dynamics behind a circular cylinder, in which sound is used as the primary feedback signal for flow regulation. The approach integrates deep reinforcement learning (DRL) with synthetic jet actuation, using acoustic measurements acquired from a downstream hydrophone array to inform control decisions in real time. Unlike conventional flow control strategies that rely on velocity or pressure field sensing, the proposed method establishes a direct link between far-field acoustic emissions and near-field actuation. Within this formulation, the DRL agent learns control policies that exploit acoustic signatures of vortex shedding to modulate synthetic jet actuation on the cylinder surface. The resulting control suppresses coherent wake structures and mitigates flow-induced disturbances. Quantitative results show reductions of up to 9.5% in radiated noise and 23.8% in drag under the tested conditions, accompanied by a marked attenuation of wake oscillations, for a DFG 2D benchmark flow with Reynolds number 100. These findings demonstrate that acoustic sensing alone can provide sufficient information for effective closed-loop flow control and highlight its potential as a non-intrusive feedback modality for coupled aerodynamic and aeroacoustic optimization in bluff-body flows. The codes for the algorithm can be found here: https://github.com/Siddharth-Rout/FlowControlDRL.
I Introduction
@@
Flow control has always been one of the most anticipated engineering problems due to its ubiquitous applicability. From the suppression of flow oscillation in open cavities Cattafesta et al. (2008) to the construction of hybrid rocket motors Tan et al. (2023), flow control has been used as an indicator of how technology has developed to counter the stochasticity of nature. Throughout the last few decades, the number of paradigms for flow control has been increasing more and more. Applications of flow control to air vehicle systems, including fixed wing airfoils, turbomachinery, combustion, aeroacoustics, vehicle propulsion integration, and rotorcraft. Flow control methods can be categorized into Active Flow Control (AFC) and Passive Flow Control (PFC). There are many innovative applications established in various industriesJoslin and Miller (2009); Wang and Feng (2018). PFC have a constant control law that is consistent with time and do not get any feedback on how well the controller performs, such as having changes to aerodynamic shapes or textures. The passive methods include Gurney flap, vortex generator, bump, cavity, roughness, small disturbance, bleed, splitter plate, polymer, and biomimetic techniquesRose et al. (2021); Othman et al. (2023). Some examples are leading-edge serrations, riblets, corrugated airfoils and lubricated skins. Most of these are widely implemented in aircrafts to delay flow separation and increase lift to drag ratio. Winglets are nowadays found to be used to reduce tip vortex formation to reduce drag. However, these control strategies are limited as the control can not be manipulated temporally based on feedback or requirements. For instance, what if the PFCs act adversely? So, AFC is a good way out as it can take in feedback from the state and actuate the controller intelligently. The active methods include oscillation and flow perturbation, acoustic excitation, jet, synthetic jet, plasma actuator, and Lorentz force. Many interesting AFCs developed in past decades. For example, installation of the synthetic jet to change the vortex shedding patternFeng and Wang (2014); utilizing wavelength actuators to attenuate turbulenceBhattacharya and Gregory (2018); and studying the effects of acoustic excitation on vortex shedding Fujisawa et al. (2004). Among those numerous methods, the application of blowing-suction of velocity jets stands out as one of the most practical and widely recognized, evidenced by NASA’s experiment on the Boeing 757 with jet actuators incorporated in the vertical stabilizer to reduce drag and improve the overall performance of the plane 28; 5.
@@
Why specifically controlling flow past a cylinder? Flow-induced forces play determinant roles in the life and safety of structures as well. Oscillations in the flow cause fatigue, enhance defects, aeroelastic flutters, and decrease the factor of safety of structures. Falling of the famous Tacoma Narrows Bridge is a popular case of structural failure due to similar causesArioli and Gazzola (2015). Tall buildings like Taipei 101, Burj Khalifa etc. have to be designed to be able to face fast windsTuan and Shang (2014); Gu et al. (2022). Passive techniques developed in the past few decades are still very promising due to the ease of utilization in industryOwen et al. (2001); Baek and Karniadakis (2009); Law and Jaiman (2017). AFCs past cylinders are excellent toy problems to demonstrate concepts. The oppositely placed suction and blowing around cylinder became popular active control strategy in 2000sKim and Choi (2005); Dong et al. (2008); Wang et al. (2016). To improve the quality nonlinear control algorithmsMao et al. (2015) and eigensystem-realization based reduced order model for suppression of wakesYao and Jaiman (2017) are popular. However, these deterministic control algorithms like proportional–integral–derivative(PID) controllers often require approximation of state space and calculation of the transfer function to actuate AFCs is expensive yet inaccurate and non-generalizable as the transfer function is case dependant. Hence, data-driven model-free methods like reinforcement learning (RL) for AFC are well appreciated as they are generalizable.
@@
In the past few years, there has been a surge in Deep Reinforcement Learning (DRL) based flow control techniquesYizhe et al. (2022); Yousif et al. (2023); Li and Zhang (2022); Fan et al. (2020). This work is hence based on a DRL algorithm to be able to extend the work to realistic cases like flow past marine vehicles. A few recent attempts to utilize DRL in AFC include controlling two synthetic jets of blowing/suction Rabault et al. (2019) and implementing adjoint-based partial derivative equation augmentation to DRL to solve flow simulation more efficientlyLiu and MacArt (2024). Both showed great results in controling wakes. In order to reduce vortex shedding, the two aforementioned papers both tried to decrease the drag coefficient in the simulation. However, no work is registered in which acoustic-based flow control model is used to reduce wakes formation and control flow dynamics. Unlike many existing DRL-based active flow control studies that rely primarily on velocity-field observations or force coefficients as feedback variables, the present work investigates acoustically informed pressure-feedback measurements as the principal control signal for adaptive wake stabilization. The study therefore focuses on the interaction between wake-induced hydrodynamic pressure fluctuations, aeroacoustic response, and DRL-guided synthetic jet actuation.
@@
Deep Q-Nework (DQN) is a branch of DRL that involves calculating the Q values of each step the model takes and, much like other ML algorithms, learning to maximize the returned reward from the environment. DQN seems to function well in more abstract tasks and therefore is widely used in robotics and has achieved great success in this field. For instance, Fernandez-Fernandez et al. study the application of DQN to human-like sketching performed by robotsFernandez-Fernandez et al. (2023). In the context of of controls and optimization in fluid dynamics morphing airfoils and shape optimization using DQN is also evidentRout and Lin (2022). Despite the level of sophistication of the model, it is not very widely used in AFC modelling. In addition, manipulation of other properties apart from drag calculation is not common in the overall scientific conversation in AFC. According to Klapwijk et al., turbulence in the fluid flows is the source of sound generation in the system. The article explores noise levels when the turbulent flow is increasedKlapwijk et al. (2022). Interestingly, it also claims that the noise generation mechanism is difficult to understand. In this work, it is shown that by controling vortex generated noise using deep Q-learning drag and wake amplitudes could be controlled.
@@
In the nineteenth century, wakes are identified as the major source of noise in the flow past an object. There lies the concept behind this work and it is theoretically supported. Strouhal in 1878 and Kohlrausch in 1881 independently found out about a faint sound originating from vortices, to which the latter described as ’reibungstone’ Strouhal (1878); Stefanini (1895); Rout (2023). Sir James Lighthill, in 1950s, discovered the theoretical connection between fluid flow and acoustics from the conservation laws to derive the wave equation for acoustics, called the theory Lighthill’s Acoustic Analogy Lighthill (1952). Sir Lighthill creates an experiment assuming a patch of turbulent flowing fluid surrounded by a large domain of surrounding stationary fluid. Let turbulent flow produce noise; however, the noise would transmit to the surrounding fluid at rest. By analysing and comparing the terms in the conservation equations for stationary fluid, the resulting equation could be written as a forced bidirectional wave equation. It becomes clear that the forcing term or the source term in the wave equation is what generates noise in flow. Here is the derived wave equation in Einstein’s notations,
$\frac{\partial^{2}\rho}{\partial^{2}t}-c_{o}^{2}{\nabla^{2}\rho}=\frac{\partial^{2}{T_{ij}}}{{\partial x_{i}}{\partial x_{j}}}$ (1)
${T_{ij}}={\rho}{v_{i}v_{j}}-{\sigma_{ij}}+(p-c_{o}^{2}\rho){\delta_{ij}}$ (2)
@@
where T is called Lighthill’s turbulence stress tensor and has three components or three sources for noise generation. ${\rho}{v_{i}v_{j}}$ is the convection of momentum fluctuation, $\mathbf{\sigma}$ is the viscous stress tensor and $(p-c_{o}^{2}\rho)\delta_{ij}$ is the difference in exact pressure $p$ and approximated thermodynamic pressure, $c_{o}^{2}\rho$. $\rho$} is the density, $c_{o}$ is the speed of sound, $t$ is the time dimension, x is the spatial dimension and v is the velocity. This equation quantifies sound sources from flowing fluid where it takes care of thermodynamics jumps, turbulent fluctuations, and viscous dissipation. Now the fact is if the stress tensor is non-zero sound is formed and with the origin of wakes, vortex stretching starts into action and even with laminar vortex street sound is produced. This sound is often considered tonal while a broadband noise is generated in case of turbulent flows. Various further works have demonstrated this Inoue and Hatakeyama (2002); Kumar et al. (2023); Liow et al. (2006).
@@
In this paper, due to the limitation of computational resources, the simulations are conducted using incompressible flow. This is based on this simple assumption that incompressible flows still produce acoustic-like pressure fluctuations, though incompressible flow solvers assume constant density. This is because the Lighthill’s stress tensor (T) is non-zero. The speed of sound in such cases tends to infinity. As such speed of sound being much slower than the speed of fluid, Mach number << 1, the nearfield observations are barely affected Crighton (1993); Meister et al. (2002); Ask et al. (2003); Layton and Novotnỳ (2009). We assume these pressure fluctuations as approximate acoustic noise. We essentially use a proxy acoustic metric derived from hydrodynamic pressure fluctuations. So, wakes make noise like pressure fluctuation and hence definitely if louder is the noise in a flow then stronger are the wake vorticities. Though this logic is very much valid but it has not been used for control or analysis rather flow states like pressure and velocity fields are considered as direct measurements and sometimes vorticity field is used as a direct measure of rotational energy in a flow. In this research, we aim to minimize the wake formation in the flow and hence lower the specially calculated effective sound pressure levels (SPLs) created by the vortices in the flow past a stationary circular cylinder using the flow-generated sound. Source of sound is a better and easier way to control vorticity in a flow as vorticity is the source of generated noise, hence being a more logical measure for such flow control problems. Furthermore, we will explore the effect of lowering SPLs on the oscillating drag experienced by the cylinder. As for the active control algorithm aspect of the research, DQN based reinforcement learning is used to control the blowing-suction of two synthetic jets on opposite ends of the cylinder perpendicular to flow.
@@
II Deep Q-Learning
Constructed on the Markov Decision Process in which the quality of action at a particular state is learned based on the reward due to the action, Q-Learning, has been a very popular early reinforcement learningWatkins (1989). Conceptually, the action at a particular state is independent of the historical state following Markov probability. However, the rewards are learned from the cumulated score of rewards in an episode of the control process. The convergence of the optimality control problem using the Bellman Equation under stochastic updates was proved soon afterWatkins and Dayan (1992). The limitation of Q-Learning is the finite nature of the map between state to best action, which is called Q-table. Deep neural networks as excellent maps in Q-Learning replace the Q-table to get called Deep Q-Network (DQN) and the algorithm is called Deep Q-Learning. It is a breakthrough reinforcement learning algorithm since it has been able to match human-level control of console to Atari gamesMnih et al. (2013). DQN allows mapping to conditional non-linear control algorithms. The second benefit is the possibility of being trained in an infinite and continuous environment state space.
Figure 1: Interaction between DQN agent and Environment in Markov Decision Process.
@@ II Deep Q-Learning
$\mathbf{V}(\mathbf{s})=max(\mathbf{R}(\mathbf{s},\mathbf{a})+\gamma\mathbf{V}(\mathbf{s^{\prime}}))$ (3)
@@
Figure 2: Diagram of flow past cylinder setup, the jets are positioned at the yellow dots shown on the top and bottom surface of the cylinder.
@@
Regarding the construction of the neural network of the DQN Agent, four fully connected layers are made between the input and output layers. The input layer takes in the calculated SPL and the output layer gives out the two jet velocities. Each layer consists of 50 neural nodes which let 7902 learnable parameters. ReLU is the nonlinear activation function used at each node, which essentially acts as switch, hence making DQN a multilayer nonlinear switch. Adam optimizerKingma and Ba (2014) in PyTorch library is applied to maximize rewards returned in each episode. More information about the utilization of the DQN agent in different tasks will be provided in the following section. The purpose of this task is to minimize the sound pressure levels (SPL) created by the wakes past the cylinder. There are multiple specific setup details in order to obtain the SPL reduction.
III Problem introduction and setup
The computational setup is built on DOLFINx, which is a high-performance solver of partial differential equations written in C++ for backend integration with legacy FEniCSx(version 2019.1.0) and python for interfaceLogg and Wells (2010); Logg et al. (2012). The project allows the use of the standard benchmark case "Flow past a cylinder (DFG 2D-3 benchmark)”, as a simulation framework to further develop the research based on Langtangen and Logg (2016); Schäfer1996; John (2004). The setup includes a horizontal rectangle with a height of 0.41m and a length of 2.2m, and the bottom left corner of the rectangle is at coordinate (0, 0). The obstacle is a circular-based cylinder with a radius of 0.05m centered at coordinate (0.2, 0.2). As the flow develops its oscillation, though laminar, the obstacle will experience a drag force, $C_{D}$, which can be determined using the formula:
$C_{D}=\frac{2}{\rho U^{2}_{mean}L}\int_{\partial\Omega_{S}}\{\rho\nu\mathbf{n.}\mathbf{\nabla}u_{t_{s}}(t){n_{y}}-p(t){n_{x}}\}ds,$ (4)